# Reducing the Scope of Language Models

### v2

@@ Abstract
We now deploy language models in a wide variety of user-facing applications. Typically, these deployments have some specific purpose, like answering questions about documentation or acting as coding assistants, but they require general language understanding. Under these circumstances these models should not be able to answer irrelevant requests such as, poetry generation or questions about physics, etc. Instead we would like language models to only answer to queries corresponding to desired behavior and refuse all other requests, which we refer to as scoping. We conduct a comprehensive empirical evaluation of potential methods from prompting to fine-tuning to preference learning to a recently proposed method for general alignment called Circuit Breakers (CB). Across three families of language models and a broad variety of tasks, we show that it is possible to scope language models. We examine scoping for multiple topics, and fine-grained topics. We ablate diversity of irrelevant queries, layer different techniques, conduct adversarial evaluations and more. Among other results, we find that, when diverse examples of irrelevant queries are available, simple supervised fine-tuning produces the best results, but when such diversity is low, Circuit Breakers perform quite well. One can often get the benefits of both methods by layering them in succession. We intend our study to serve as a practitioner’s guide to scoping language models.
@@ 1 Introduction
In the past few years, large language models have exploded into the popular conscience. One major recent addition is the “alignment” process through Reinforcement Learning with Human Feedback (RLHF) (Christiano et al., 2017; Ouyang et al., 2022), which has made the current generation of language models much less likely to emit toxic content than previous generations (Wolf et al., 2017), and thus much more acceptable for general use. As a result, many businesses and individuals feel more comfortable using these technologies than they would be in the past.
@@
While we have generally capable language models that refuse to answer toxic or dangerous queries, it is still difficult to deploy these language models. Even though they may not emit toxic content, they will still happily answer any question, irrelevant or not. This becomes a problem when we wish to use them in specific contexts: e.g. shopping bots currently give coding advice [^1] https://shorturl.at/qf3FA or answer other questions, [^2] https://shorturl.at/R0JDv while assistive copilots can be taken off course by prompt injections. [^3] https://shorturl.at/yyU6P Thus, there is still a need to scope language models for these specific uses.
@@
Currently, language models can be scoped through two-stage approaches like relevance classifiers, or system prompting, but these options are brittle (Chao et al., 2023; Mehrotra et al., 2023; Zeng et al., 2024; Wei et al., 2023) and easy to circumvent. We shed further light on this problem, conducing a comprehensive empirical study on scoping language models to specific tasks. We apply existing methods to this problem, including system prompting, supervised fine-tuning, preference learning (Rafailov et al., 2024), probing, and a recently-introduced method called Circuit Breakers (Zou et al., 2024). We scope language models more broadly for multiple tasks, and more finely for specific niche tasks. We ablate over diversity of training sets, language model size, adversarial prompting and more. Our specific contributions include:
@@
  • We show it is possible to scope language models, even for multiple and fine-grained tasks
  • We find when training data is diverse, supervised fine-tuning is best
  • When data diversity is low, Circuit Breakers (Zou et al., 2024) is better
  • We show it is possible to layer these two, often preserving the best performance
@@ 2 Related Work
Figure 1: We study the ability to scope of language models to specific topics. We assume access to a set of positive (accept) queries and negative (reject) queries, where the accept queries correspond to a relatively narrow domain. We examine how well different methods cause the language model to accept only the positive examples, while rejecting all other examples, including out-of-distribution requests that weren’t seen during training.
Aligning Language Models: The advent of the current era of language models has been marked by a process of aligning language models so that generations are more helpful, and safer for deployment (Ouyang et al., 2022; Bai et al., 2022a). The primary way this is accomplished is through reinforcement learning with human feedback (RLHF) (Christiano et al., 2017) which was first proposed in robotic simulation tasks. RLHF proceeds by collecting preference pairs of completions, and training a reward model from human judgments on those preference pairs, then performing reinforcement learning with the language model against that reward model. From tasks in simulation, it was developed in language (Stiennon et al., 2020), until it reached its current state. Other works have removed the human aspect of human feedback, allowing for synthetic feedback from models (Bai et al., 2022b; Sudalairaj et al., 2024). Lately, Rafailov et al. (2024) have removed the need for a reward model, making for a stabler and simpler objective function without many of the complexities of RL training. A budding line of work also explores aligning not just to a single reward model, but preferences of many different individual users (Chakraborty et al., 2024; Lee et al., 2024). All of these methods focus on some general notion of alignment, without considering the specific task, unlike our work.
Adapting for Specific Purposes: Typically after pretraining, language models go through an instruction fine-tuning stage, where they gain the ability to follow instructions (Mishra et al., 2022; Ouyang et al., 2022; Wei et al., 2022). After this, they proceed through an alignment phase as discussed above, usually to avoid harmful behavior (Bai et al., 2022a). It is possible to adapt language models for specific purposes simply with a system message (Touvron et al., 2023), but many examples of black-box adversarial attacks (Chao et al., 2023; Anil et al., 2024; Wei et al., 2023; Zeng et al., 2024) demonstrate it is difficult only to rely on the system prompt for such control. Wallace et al. (2024) propose finetuning with different levels of priority, similar to Zhang et al. (2023b), but these works focus primarily on general safety and not the task. These ideas are based on the fact that current language models can often be distracted by irrelevant context (Shi et al., 2023; Yoran et al., 2024). Thus, it seems important to finetune the language model if we want it to be deployed to a particular domain. For domains where there is sufficient data, we may also pretrain and fix the language model’s purpose ahead of time (Beltagy et al., 2019; Wu et al., 2023; Li et al., 2023) or continue pretraining from a base language model (Gururangan et al., 2020). It is an open question however whether finetuning retains the robustness capabilities, or if it is similarly as brittle as system prompting for out-of-distribution questions.
Refusal in Language Models: As our work deals with scoping models to refuse irrelevant queries, we review refusal. More detail is available in a comprehensive survey by Wen et al. (2024). One common case to train for refusal is when the answer is unknown or the model is unconfident (Zhang et al., 2023a; Cao, 2023; Xu et al., 2024). Another is for unsafe inputs (Varshney et al., 2023; Zhang et al., 2023b; Wallace et al., 2024). Supervised fine-tuning (SFT) to reject unsafe prompts can still lead to unsafe behavior, though parameter efficient methods like LoRA (Hu et al., 2022) have better tradeoffs (Brahman et al., 2024). Both Brahman et al. (2024) and Cheng et al. (2024) take an approach to refusal using SFT and DPO, which we will adapt to our case. Other methods to induce refusal may be prompt-based (Xie et al., 2023; Zhang et al., 2024) or based on probing model representations (Kadavath et al., 2022; Slobodkin et al., 2023). Zou et al. (2024) design a method that conditionally rejects unsafe inputs based on orthogonalizing internal representations that we will adapt for our study. Though these methods lay out a set of techniques to explore for our task, all of them are oriented toward general alignment qualities like safety, as opposed to specific tasks that we will explore.
3 Experimental Setup
3.1 Scoping for Specific Tasks
We would like to scope language models to provide completions to relevant tasks, and reject queries corresponding to irrelevant tasks. In particular, we assume we are given a set of “accept” queries $\{x_{a}|x_{a}\sim\mathcal{A}_{k}\}$, where $\{\mathcal{A}_{k}\}$ is a set of “accept” tasks, and a set of “reject” queries $\{x_{r}|x_{r}\sim\mathcal{R}_{k}\}$ where $\{\mathcal{R}_{k}\}$ is a set of ”reject” tasks. We are given a language model $f_{\theta}:x\mapsto y$ which predicts completion $y$ from input $x$, with parameters $\theta$; a classifier $g:y\mapsto c\in\{0,1\}$ which decides whether a completion is accepted (0) or rejected (1) by the language model. We would like to compute an update $\Delta$ such that we minimize $\mathop{\mathbb{E}}_{x_{a}}g(f_{\theta+\Delta}(x_{a}))$ and maximize $\mathop{\mathbb{E}}_{x_{r}}g(f_{\theta+\Delta}(x_{r}))$. Thus we want ‘accept’ queries to be accepted and ‘reject’ queries to be rejected.
As an additional goal, we would like performance on the accept tasks not to degrade. Given a scoring function $h:(x,y)\mapsto s\in[0,1]$ which scores the completion on task performance where 1 is best, we would also like to maximize $\mathop{\mathbb{E}}_{x_{a}}h(x_{a},f_{\theta+\Delta}(x_{a}))$.
3.2 Datasets
We conduct many experiments with different mixtures of accept and reject queries. In order to standardize the format, we draw prompts from Super-NaturalInstructions (SNI) (Wang et al., 2022). SNI is a meta-dataset composed of many different “tasks”, sometimes with multiple tasks per dataset, for example generating questions from passages for a reading comprehension dataset, or generating answers to provided questions from the same reading comprehension dataset. Each task, specified by a task instruction, comes with a collection of examples. We use SNI as it is publicly available, and contains a broad range of complex tasks which current language models should be able to perform. To get our training datasets, we first manually select a set of tasks that are straightforward to automatically evaluate, leaving out many more subjective tasks that may require a human reader. We then group those tasks that we select by category provided from SNI. Details and statistics on categories are provided in Table 1.
@@ 3.2 Datasets
Sentiment Analysis (SA) Predicing if a movie review is positive or not 8 10 31248
@@
Tex Completion (TC) Filling in the blanks in a transcript 3 3 10515
@@
GSM8k (Cobbe et al., 2021) Answering simple math word problems 1 1 5978
Alpaca (Taori et al., 2023) General requests like writing a recipe for lunch 1 - 18793
@@
Each of these categories contains multiple datasets, so the distribution for each task is quite broad. We will also combine multiple tasks in the accept or reject set. For all experiments, we always evenly split the training data for accept/reject set between all tasks. We reserve at least 20% of the prompts as a validation set that are not seen during training. Where not specified, we use 2048 prompts for the accept set, and 2048 prompts for the reject set. We evaluate Sentiment Analysis and Toxic Language Detection with accuracy (the classes are mostly balanced), while for other tasks we use a standard metric for generation, Rouge-L (Lin, 2004), between the generation and ground truth completion as a proxy for performance (Accept Score). Our goal is mostly to study rejection behavior, so a rough performance proxy is all we need.
@@
Regardless of the accept and reject sets used for training, we will evaluate all of the categories given in Table 1, then present averages of rejection rates in the accept (Accept), all sets in in-distribution reject (ID Reject) and all sets in out of distribution reject (OOD Reject). Note if the training set consists of SA in Accept, and S in TLD Reject, OOD Reject will contain the 8 other categories. Tasks above the divider in Table 1 will be used in different experiments for both training and evaluation, while tasks below the divider will only be used for OOD evaluation. As stated previously, ideally we would like to have 0 rejection on Accept, and 100% rejection on ID Reject and OOD Reject.
3.3 Methods
@@
For all methods that require training the language model, we use LoRA (Hu et al., 2022) with default rank 16, $\alpha=16$ and dropout of 0.05. We use Adam (Kingma, 2014) without any regularization and tune learning rates (see Appendix B).
@@
System Prompting (Sys.): The simplest method to scope language models is simply to instruct them to refuse irrelevant prompts. For example, for SA the system is: You are an assistant who only answers request related to Sentiment Analysis. For all other requests you respond “I cannot answer that.” With multiple accept categories, we comma separate the category names (e.g. “…related to Sentiment Analysis, Text Completion and Summarization…”). This system prompt is prepended to all instructions at evaluation time. In addition, all other methods also use the system prompt both at training and evaluation time. This is similar to methods proposed by Xie et al. (2023); Zhang et al. (2024).
@@
Supervised Fine-Tuning (SFT): Supervised Fine-Tuning (SFT) consist of tuning the language model to produce particular outputs. For the accept tasks the completions $y_{a}$ are the groundtruth completions provided by the dataset. For the reject tasks, the completions $y_{r}$ are always “ cannot answer that.”. As the finetuning dataset can be quite small, loss is only computed on the completions so as to avoid overfitting to the small set of instructions, agreeing with common practice (Mishra et al., 2022; Ouyang et al., 2022; Wei et al., 2022). We tune learning rate and step budget for SFT. This is similar approach to Brahman et al. (2024); Cheng et al. (2024).
@@
Direct Preference Optimization (DPO): Given its role in post-training we want to explore a preference learning method. We choose to experiment on Direct Preference Optimization (DPO) (Rafailov et al., 2024), as it does not require an additional reward model. DPO requires pairs of preference data, so for accept queries we provide the dataset completion as preferred, and the completion “I cannot answer that.” as rejected. For reject queries we do the reverse, preferring “I cannot answer that.” over the ground truth completion. For DPO we tune learning rate, step budget, and the loss weighting term regularizing the KL divergence from the base model predictions. This is similar to Brahman et al. (2024); Cheng et al. (2024).
@@
Probing Classifier (Probe): Probes of representations are a common method to accomplish tasks as they base predictions on the internal state of the language model (Conneau et al., 2018; Tenney et al., 2019; Zou et al., 2023a). Previous work on Circuit Breakers (Zou et al., 2024) showed that probing representations was competitive for detecting dangerous language. However, they only designed probes on a single layer of a language model. Here we design a stronger probe. Once an instruction is fed to the frozen language model, we first remove the first position as it is quite anomalous due to large magnitude (Xiao et al., 2024), then we average all positions per layer and normalize the average vector to norm 1 so as to match norms between layers. Finally we concatenate the average vectors from each layer and feed that as input to a 2-layer MLP with width 256 which makes a binary classification decision on whether to accept or reject. Only the MLP layers are trained, and we tune the learning rate and step budget. This is justified by prior choices in Kadavath et al. (2022); Slobodkin et al. (2023); Zou et al. (2024).
@@
Circuit Breakers (CB): Zou et al. (2024) first introduce a method they call Circuit Breakers (CB) for accepting normal queries while rejecting dangerous ones. We repurpose their method for this task. Essentially given a function rep which extracts the representations of a language model at particular layers, they design an optimization objective with two components: $\mathcal{L}_{a}(x_{a},\Delta)=\lVert\text{rep}(f_{\theta}(x_{a}))-\text{rep}(f_{\theta+\Delta}(x_{a}))\rVert_{2}^{2}$ and $\mathcal{L}_{r}(x_{r},\Delta)=\max\{0,\cos(\text{rep}(f_{\theta}(x_{r})),\text{rep}(f_{\theta+\Delta}(x_{r}))\}$. The total loss is $\mathcal{L}=\alpha(t)\mathcal{L}_{a}+\beta(t)\mathcal{L}_{r}$ where the two components of the loss are scheduled over time.
@@
3.4 Detecting rejection
Though a strong language model judge (Zheng et al., 2023) may seem like a good choice for judging rejection, we followed prior work in first experimenting with simpler heuristics (Zou et al., 2023b; 2024) based on string matching and the rejection behavior of Circuit Breakers. We found these heuristics to very strong, and much less expensive than running a frontier judge for the many experiments in this paper. We also experimented with a smaller hosted language model judge, but found its performance much poorer than the heuristics. More details on evaluation are available in Appendix B.
4 Experiments
All experiments contain of evaluations of task performance (Accept Score) on the accept set (which should be high), rejection rate on the in-distribution accept (Accept) set (which should be low) as well as rejection rate on the in-distribution reject set (ID Reject) and out of distribution data (OOD Reject) (which should be high). We describe experiments in broad strokes, and defer precise details on hyperparameters to Appendix B. We begin by presenting results across a variety of language models (Mistral-7B-Instruct-v0.2, granite-7b-instruct, Llama-3.1-8B-Instruct(Jiang et al., 2023; Sudalairaj et al., 2024; Grattafiori et al., 2024). We explore scale on the Llama-3.2-instruct family, and for other results choose Mistral-7b-instruct as we do not have the compute or space to run every experiment on every model.
4.1 Scoping language models
We start with the basic question: is it possible to scope language models as we set out to? We explore each method across a variety of models using Sentiment Analysis (SA) as our accept task in Figure 2, where we see that although system prompting is insufficient, a broad variety of different methods are successful to different extents with different models.
@@ 4.2 Rejection set diversity
One of the most critical questions when attempting to restrict the generations of language models is what data might be necessary to do so. If models overfit to a particular data distribution, then it may be difficult to reject requests that were not specified in the training distribution. Thus, here we ask: how much data diversity is necessary in the rejection set to robustly scope models? If very little diversity is needed, and rejection extends to OOD requests, then adapting models to new deployments becomes quite inexpensive. In Appendix Figure 3 we study Sentiment Analysis across different models with more and more diverse rejections sets consisting of an increasing number of tasks, and defer results on additional tasks with Mistral-7b-instruct to Appendix A.1.
4.3 Accepting multiple tasks
Here we ask: is it possible to still reject tasks well when there are multiple tasks in the rejection set? Such a setting is natural as most language models will have a few different specific uses, like a programming bot that can write code and also answer questions about documentation. We demonstrate results on Mistral-7b-Instruct-v0.2 in Table 2 with two choices of accepts sets: Classification and Generation (SA,S,TLD,SC,TC,DG) and Math and Programming (PE,GSM8k). We use the opposite set as the ID reject set in this case.

### v3

@@ Abstract
Large language models (LLMs) are deployed in a wide variety of user-facing applications. Typically, these deployments have some specific purpose, like answering questions grounded on documentation or acting as coding assistants, but they require general language understanding. In such deployments, LLMs should respond only to queries that align with the intended purpose and reject all other requests, such as generating poetry or answering questions about physics, a task we refer to as ‘scoping’. We conduct a comprehensive empirical evaluation of various methods, ranging from prompting, fine-tuning to preference learning and the recently proposed general alignment technique known as Circuit Breakers (CB). Across three families of language models and a broad variety of tasks, we show that it is possible to scope language models. We examine scoping for multiple topics, and fine-grained topics. We ablate diversity of irrelevant queries, layer different techniques, conduct adversarial evaluations and more. Among other results, we find that when diverse examples of irrelevant queries are available, simple supervised fine-tuning produces the best results, but when such diversity is low, Circuit Breakers perform quite well. One can often get the benefits of both methods by layering them in succession. We intend our study to serve as a practitioner’s guide to scoping LLMs.
Codehttps://github.com/IBM/llm-scoping
@@ 1 Introduction
In recent years, large language models have surged into public awareness. One major recent addition is the “alignment” process through Reinforcement Learning with Human Feedback (RLHF) (christiano2017deep; ouyang2022training), which has made the current generation of language models much less likely to emit toxic content than previous generations (wolf2017we), and thus much more acceptable for general use. As a result, many businesses and individuals now feel more comfortable using these technologies than they did in the past.
@@
Although we have generally capable language models that can refuse to answer toxic or dangerous queries, deploying them still remains challenging. Even if they avoid producing harmful content, they often respond to any question, relevant or not, without discernment. This becomes a problem when we wish to use them in specific contexts: e.g. shopping bots currently give coding advice [^1] https://shorturl.at/qf3FA or answer other questions, [^2] https://www.forbes.com/sites/lesliekatz/2024/07/13/amazon-ai-shopping-assistant-rufus-answers-non-shopping-questions-too/ while assistive copilots can be taken off course by prompt injections. [^3] https://oecd.ai/en/incidents/2025-10-08-0fbf Thus, there is still a need to scope language models for these specific uses.
Figure 1: We study the ability to scope language models to specific topics. We assume access to a set of relevant (accept) queries and irrelevant (reject) queries, where the accept queries correspond to a relatively narrow domain. We examine how well different methods cause the language model to accept only the relevant examples, while rejecting all other examples, including out-of-distribution requests that weren’t seen during training.
We define LLM scoping as a conditional generation task in which a language model must: (i) identify whether an input query falls within a relevant domain, (ii) reject irrelevant queries, and (iii) maintain high-quality generation for relevant queries. This contrasts with traditional text classification, which maps inputs to discrete labels without requiring natural language generation.
Let $\mathcal{Q}$ denote the set of all possible natural language queries. Let $\mathcal{D}_{\text{rel}}\subset\mathcal{Q}$ denote the subset of relevant or in-domain (should be accepted) queries, and let $\mathcal{D}_{\text{irr}}=\mathcal{Q}\setminus\mathcal{D}_{\text{rel}}$ denote the set of irrelevant or out-of-domain (should be rejected) queries. Let $\mathcal{R}$ be the space of valid natural language responses, and let $\bot\notin\mathcal{R}$ be a special token representing rejection.
The scoping is a function:
$f_{\theta}:\mathcal{Q}\rightarrow\mathcal{R}\cup\{\bot\}$
such that:
$f_{\theta}(q)$ $\in\mathcal{R},\quad\text{if }q\in\mathcal{D}_{\text{rel}},$
$f_{\theta}(q)$ $=\bot,\quad\text{if }q\in\mathcal{D}_{\text{irr}}.$
In contrast, text classification models are not required to preserve generation quality, whereas in LLM scoping, degrading the model’s performance on $\mathcal{D}_{\text{rel}}$ constitutes a failure of the task. Further, unlike the traditional LLM refusal task in the context of safety and alignment, which typically corresponds to a much smaller reject set than accept set, the scoping tasks consider the opposite.
@@
Currently, LLMs can be scoped through two-stage approaches like relevance classifiers, or system prompting and then text generation, but these options are brittle (chao2023jailbreaking; mehrotra2023tree; zeng2024johnny; wei2024jailbroken) and easy to circumvent. We shed further light on this problem, conducting a comprehensive empirical study on scoping language models to specific tasks. We apply existing methods to this problem, including system prompting, supervised fine-tuning, preference learning (rafailov2024direct), probing, and a recently-introduced method called Circuit Breakers (zou2024improving). We scope language models more broadly for multiple tasks, and more finely for specific niche tasks. We ablate over diversity of training sets, language model size, adversarial prompting and more. Our specific contributions include:
@@
  • We show that it is possible to scope language models, even for multiple and fine-grained tasks
  • We find that when training data exhibits high diversity, supervised fine-tuning yields the best performance
  • Conversely, in settings with low data diversity, the Circuit Breakers method (zou2024improving) provides superior results
  • Finally, we show it is possible to layer these two, often preserving the best performance
@@ 2 Related Work
Sentiment Analysis (SA) Predicing if a movie review is relevant or not 8 10 31248
@@
Text Completion (TC) Filling in the blanks in a transcript 3 3 10515
@@
GSM8k (cobbe2021training) Answering simple math word problems 1 1 5978
Alpaca (taori2023alpaca) General requests like writing a recipe for lunch 1 - 18793
@@
Aligning Language Models: The advent of the current era of language models has been marked by a process of aligning language models so that generations are more helpful, and safer for deployment (ouyang2022training; bai2022training). The primary way this is accomplished is through reinforcement learning with human feedback (RLHF) (christiano2017deep) which was first proposed in robotic simulation tasks. RLHF proceeds by collecting preference pairs of completions, and training a reward model from human judgments on those preference pairs, then performing reinforcement learning with the language model against that reward model. From tasks in simulation, it was developed in language (stiennon2020learning), until it reached its current state. Other works have removed the human aspect of human feedback, allowing for synthetic feedback from models (bai2022constitutional; sudalairaj2024lab). Lately, rafailov2024direct have removed the need for a reward model, making for a stabler and simpler objective function without many of the complexities of RL training. A budding line of work also explores aligning not just to a single reward model, but preferences of many different individual users (chakraborty2024maxmin; lee2024aligning). All of these methods focus on some general notion of alignment, without considering the specific task, unlike our work.
Adapting for Specific Purposes: Typically after pretraining, language models go through an instruction fine-tuning stage, where they gain the ability to follow instructions (mishra2022cross; ouyang2022training; wei2022finetuned). After this, they proceed through an alignment phase as discussed above, usually to avoid harmful behavior (bai2022training). It is possible to adapt language models for specific purposes simply with a system message (touvron2023llama), but many examples of black-box adversarial attacks (chao2023jailbreaking; anil2024many; wei2024jailbroken; zeng2024johnny) demonstrate it is difficult only to rely on the system prompt for such control. wallace2024instruction propose finetuning with different levels of priority, similar to zhang2023defending, but these works focus primarily on general safety and not the task. These ideas are based on the fact that current language models can often be distracted by irrelevant context (shi2023large; yoran2024making). Thus, it seems important to finetune the language model if we want it to be deployed to a particular domain. For domains where there is sufficient data, we may also pretrain and fix the language model’s purpose ahead of time (beltagy2019scibert; wu2023bloomberggpt; li2023starcoder) or continue pretraining from a base language model (gururangan2020don). It is an open question however whether finetuning retains the robustness capabilities, or if it is similarly as brittle as system prompting for out-of-distribution questions.
Refusal in Language Models: As our work deals with scoping models to refuse irrelevant queries, we review refusal. More detail is available in a comprehensive survey by wen2024art. One common case to train for refusal is when the answer is unknown or the model is unconfident (zhang2023r; cao2023learn; xu2024rejection). Another is for unsafe inputs (varshney2023art; zhang2023defending; wallace2024instruction). Supervised fine-tuning (SFT) to reject unsafe prompts can still lead to unsafe behavior, though parameter efficient methods like LoRA (hu2022lora) have better tradeoffs (brahman2024art). Both brahman2024art and cheng2024can take an approach to refusal using SFT and DPO, which we will adapt to our case. Other methods to induce refusal may be prompt-based (xie2023defending; zhang2024intention) or based on probing model representations (kadavath2022language; slobodkin2023curious). zou2024improving design a method that conditionally rejects unsafe inputs based on orthogonalizing internal representations that we will adapt for our study. Though these methods lay out a set of techniques to explore for our task, all of them are oriented toward general alignment qualities like safety, as opposed to specific tasks that we will explore.
3 Experimental Setup
We would like to scope language models to provide completions to relevant tasks, and reject queries corresponding to irrelevant tasks. In particular, we assume we are given a set of “relevant” or “accept” queries $\{q_{rel}|q_{rel}\sim\mathcal{D}_{rel}\}$, where $\{\mathcal{D}_{rel}\}$ is a set of accepted tasks, and a set of “irrelevant” queries $\{q_{irr}|q_{irr}\sim\mathcal{D}_{irr}\}$ where $\{\mathcal{D}_{irr}\}$ is a set of rejected tasks. We are given a language model $f_{\theta}:q\mapsto y$ which predicts completion $y$ from input $q$, with parameters $\theta$; a classifier $g:y\mapsto c\in\{0,1\}$ which decides whether a LLM completion is accepted (0) or rejected (1). We would like to compute an update $\Delta$ such that we minimize $\mathop{\mathbb{E}}_{q_{rel}}g(f_{\theta+\Delta}(q_{rel}))$ and maximize $\mathop{\mathbb{E}}_{q_{irr}}g(f_{\theta+\Delta}(q_{irr}))$. Thus we want ‘relevant’ queries to be accepted and ‘irrelevant’ queries to be rejected.
As an additional goal of scoping, we would like performance on the accept tasks not to degrade. Given a scoring function $h:(q,y)\mapsto s\in[0,1]$ which scores the completion on task performance where 1 is best, we would also like to maximize $\mathop{\mathbb{E}}_{q_{rel}}h(q_{rel},f_{\theta+\Delta}(q_{rel}))$.
Datasets & Metrics
We conduct many experiments with different mixtures of accept and reject queries. In order to standardize the format, we draw prompts from Super-NaturalInstructions (SNI) (wang2022super). SNI is a meta-dataset composed of many different “tasks”, sometimes with multiple tasks per dataset, for example generating questions from passages for a reading comprehension dataset, or generating answers to provided questions from the same reading comprehension dataset. Each task, specified by a task instruction, comes with a collection of examples. We use SNI as it is publicly available, and contains a broad range of complex tasks which current language models should be able to perform. To get our training datasets, we first manually select a set of tasks that are straightforward to automatically evaluate, leaving out many more subjective tasks that may require a human reader. We then group those tasks that we select by category provided from SNI. Details and statistics on categories are provided in Table 1.
@@ 3 Experimental Setup
Each of these categories contains multiple datasets, so the distribution for each task is quite broad. We will also combine multiple tasks in the accept or reject set. For all experiments, we always evenly split the training data for accept/reject set between all tasks. We reserve at least 20% of the prompts as a validation set that are not seen during training. Where not specified, we use 2048 prompts for the accept set, and 2048 prompts for the reject set. We evaluate Sentiment Analysis and Toxic Language Detection with accuracy (the classes are mostly balanced), while for other tasks we use a standard metric for generation, Rouge-L (lin2004rouge), between the generation and ground truth completion as a proxy for performance (Accept Score). Our goal is mostly to study rejection behavior, so a rough performance proxy is all we need.
@@
Irrespective of the specific accept and reject sets used during training, we evaluate performance across all categories listed in Table 1. We then report the average rejection rates separately for the accept set (Accept), all in-distribution reject sets (ID Reject), and all out-of-distribution reject sets (OOD Reject). Note if the training set consists of SA in Accept, and S in Reject, OOD Reject will contain the 8 other categories. Tasks above the divider in Table 1 will be used in different experiments for both training and evaluation, while tasks below the divider will only be used for OOD evaluation. As stated previously, ideally we would like to have 0 rejection on Accept, and 100% rejection on ID Reject and OOD Reject.
Methods
@@
For all methods that require training the language model, we use LoRA (hu2022lora) training with rank $16$, $\alpha=16$ and dropout of $0.05$. We use the Adam (kingma2014adam) optimizer without any regularization and tune learning rates (see Appendix B).
@@
System Prompting (Sys.): The simplest method to scope language models is simply to instruct them to refuse irrelevant prompts. For example, for SA the system prompt is: You are an assistant who only answers requests related to Sentiment Analysis. For all other requests you respond “I cannot answer that.” With multiple accept categories, we comma separate the category names (e.g. “…related to Sentiment Analysis, Text Completion and Summarization…”). This system prompt is prepended to all instructions at evaluation time. In addition, all other methods also use the system prompt both at training and evaluation time. This is similar to methods proposed by xie2023defending; zhang2024intention.
@@
Supervised Fine-Tuning (SFT): Supervised Fine-Tuning (SFT) consists of tuning the language model to produce particular outputs. For the accept tasks the completions $y_{rel}$ are the groundtruth completions provided by the dataset. For the reject tasks, the completions $y_{irr}$ are always “I cannot answer that.”. We tune learning rate and step budget for SFT. This is a similar approach to brahman2024art; cheng2024can. In experiments, the $\mathcal{L}_{gen}$ is about generation (task completion) loss between $y_{rel}$ and $f_{\theta}(q),{q\in\mathcal{D}_{\text{rel}}}$, and the $\mathcal{L}_{rej}$ is about irrelevant task rejection loss between $y_{irr}$ and $f_{\theta}(q),{q\in\mathcal{D}_{\text{irr}}}$. The total scoping loss is $\mathcal{L}_{scope}=\mathcal{L}_{gen}+\lambda\mathcal{L}_{rej}$. We use balanced scoping loss where $\lambda=1$ for loss computing. As the finetuning dataset can be quite small, loss is only computed on the completions so as to avoid overfitting to the small set of instructions, agreeing with common practice (mishra2022cross; ouyang2022training; wei2022finetuned).
@@
Direct Preference Optimization (DPO): Given its role in post-training we explore a preference learning method. We choose to experiment on Direct Preference Optimization (DPO) (rafailov2024direct), as it does not require an additional reward model. DPO requires pairs of preference data, so for accept queries we provide the dataset completion as preferred, and the completion “I cannot answer that.” as rejected. For reject queries we do the reverse, preferring “I cannot answer that.” over the ground truth completion. For DPO we tune learning rate, step budget, and the loss weighting term regularizing the KL divergence from the base model predictions. This is similar to brahman2024art; cheng2024can.
@@
Two-stage on Probing Classifier (Probe): Probes of representations are a common method to accomplish tasks as they base predictions on the internal state of the language model (conneau2018cram; tenney2019you; zou2023representation). Previous work on Circuit Breakers (zou2024improving) showed that probing representations was competitive for detecting dangerous language. However, they only designed probes on a single layer of a language model. Here we design a stronger probe. Once an instruction is fed to the frozen language model, we first remove the first position as it is quite anomalous due to large magnitude (xiao2023efficient), then we average all positions per layer and normalize the average vector to norm 1 so as to match norms between layers. Finally, we concatenate the mean-pooled representations from each layer to form a single feature vector, which is then passed through a two-layer multilayer perceptron (MLP) with a hidden dimension of 256 to perform binary classification, determining whether to accept or reject the input. Only the MLP layers are trained, and we tune the learning rate and step budget. This is justified by prior choices in kadavath2022language; slobodkin2023curious; zou2024improving. We use this probe in a two-stage setup where the probe first classifies the request and if it is determined to be an ‘accept’ task, a response generated via an LLM.
@@
Circuit Breakers (CB): (zou2024improving) first introduce a method they call Circuit Breakers (CB) for accepting normal queries while rejecting dangerous ones. We repurpose their method for this task. Essentially, given a function which extracts the representations of a language model at particular layers, they design an optimization objective with two components: $\mathcal{L}_{gen,{q\in\mathcal{D}_{\text{rel}}}}(q,\Delta)=\lVert\text{rep}(f_{\theta}(q))-\text{rep}(f_{\theta+\Delta}(q))\rVert_{2}^{2}$ and $\mathcal{L}_{rej,{q\in\mathcal{D}_{\text{irr}}}}(q,\Delta)=\max\{0,\cos(\text{rep}(f_{\theta}(q)),\text{rep}(f_{\theta+\Delta}(q))\}$. The total loss is $\mathcal{L}_{scope}=\alpha(t)\mathcal{L}_{gen}+\beta(t)\mathcal{L}_{rej}$ where the two components of the loss are scheduled over time.
@@
Detecting rejection
@@
Though a strong language model judge (zheng2023judging) may seem like a good choice for judging rejection, we followed prior work in first experimenting with simpler heuristics (zou2023universal; zou2024improving) based on string matching and the rejection behavior of Circuit Breakers. We found these heuristics to be very strong, and much less expensive than running a frontier judge for the many experiments in this paper. We also experimented with a smaller hosted language model judge, but found its performance much poorer than the heuristics. More details on evaluation are available in Appendix B.